# Bayesian estimation of cross-coupling and reflection systematics in 21cm array visibility data

### v1

@@ 6 Summary
In this work, we present statistical modelling results when applied to the reflection and cross-coupling systematics in simulations of early HERA Phase I visibilities. These systematics serve to complicate setting upper limits on the EoR 21-cm signal by spreading copies of the foregrounds into regions ideally only occupied by the EoR. We mimic this by adding systematics with randomised parameters to visibilities containing foregrounds and the 21-cm signal, where the initial, uncorrupted visibilities are sourced from the HERA Validation data (Aguirre et al., 2022), and the systematics are added with hera_sim. A Hamiltonian Monte Carlo sampler is used to constrain these systematics, subject to noise, and attempts are made to remove them and recover the true foreground + EoR signal. We consider both an autocorrelation case, as well as a short, 14.7 m east-west cross-correlation.
A notable limitation of our method is that the foreground, EoR, and systematics models cannot distinguish between their respective components when being applied to mock data, particularly in regions containing densely distributed systematics. For a particular delay, the total power can be attributed to varying combinations of power from these three components. In practice, this resulted in less effective mitigation of the systematics.
This appears to be primarily a concern for the autocorrelations, and not for cross-correlations, as the foregrounds have a slower reduction in power towards higher delays (Fig. 7), and had a higher density of systematics. For cross-correlations, only ten cross-coupling spectra are of concern. For autocorrelations, the twenty subreflections, in particular, likely makes it difficult for the sampler to disentangle the power contributions from the multiple overlapping components. Indeed, at delays above $\sim 1100\,\rm{ns}$ in Fig. 10 the only contributions to the power are the EoR, a single reflection, and a handful of cross-coupling spectra which are less densely distributed than subreflections. Here, the level of residual systematics is mostly negligible as there are fewer contributions to the total power at a particular delay, leading to more effective mitigation.
To reduce the level of residual systematics would likely require very specific priors on either the systematics, or the foreground and EoR models. Very good knowledge of the latter components is likely not realistic, especially for the EoR given the difficulty involved with its observational constraint. Alternate modelling methods could possibly improve performance, however. As opposed to eigenvectors for the foregrounds, and Fourier modes for the EoR, it is possible that templates could provide more favourable performance, i.e. visibilities with the correct shape in frequency and delay space for which only a small number of amplitudes are adjusted in order to model these components. As it is currently implemented, the eigenvectors and Fourier modes can more easily fit individual systematics peaks. Templates could be slightly more resilient to this, as they would need to fit larger features, and in turn might disregard the finer ones.
@@
In autocorrelation power spectra, there are three systematics of concern: cable reflections, cable subreflections, and cross-coupling. These are analysed in Section 5.1. When considering a single 10.73 second time integration, we are able to constrain a number of high-amplitude reflection systematics (Fig. 8), but can only mitigate down to the (relatively high) noise level (Fig. 9), with all of the cross-coupling features being below the noise level. These reflection systematics are modelled in the autocorrelations and applied to their lower-amplitude counterparts in the corresponding cross-correlations. For the same time integration, however, reflections in cross-correlations are overwhelmed by the noise. As such, there is not much to be gained from calibrating out reflections in the cross-coupling, in this case.
@@
When the model is run on visibilities with very low noise levels, however, the systematics in the autocorrelations can be mitigated such that the recovered power spectrum is, at worst, around two orders of magnitude above the true power spectrum (Fig. 10).
Using these results to subtract the reflections from the corresponding low noise cross-correlation, we are able to essentially recover the true power spectrum in Fig. 11 (top). In this low-noise case, there is minimal signal loss or residual systematics (Fig. 11, bottom). Furthermore, as a result of the low noise, the constraints placed on the systematics parameters become very strong, resulting in narrow statistical uncertainty estimates.
In both the high-noise and low-noise cases, the non-linear optimisation method of Kern et al. (2019) performs similarly to our sampler. In the former case, systematics are mitigated to the noise level, and in the latter case there are some residual systematics in the autocorrelation power spectrum, but very good recovery of the cross-correlation with minimal signal loss. It is in the cases where noise is more realistic that the value of a Bayesian approach is more evident, as it provides some measure of statistical uncertainty in the results. For lower noise levels, these estimates become narrower than is likely useful.
@@
For a single time integration, the strongest cross-coupling peaks are about as strong as the noise, eliminating any possibility of meaningful constraints being placed the systematics. Instead, some amount of noise reduction was required. Assuming a simple time integration increase from $\Delta t=10.73\,\rm{s}$ to $1000\,\rm{s}$ reduces the noise enough to allow constraints to be placed on the highest amplitude cross-coupling peaks in Fig. 12, leading to their mitigation down to the noise level in Fig. 13.
@@
We evaluate a very-low-noise case in order to evaluate the model in an ideal scenario, i.e. where the noise power spectrum is around $10\%$ that of the EoR power spectrum. This required a slightly different model setup in comparison to all other results, but we are able to effectively mitigate the cross-coupling spectra down to the true power spectrum (Fig. 14, top). Kern et al. (2019)’s SVD method performs similarly in this low-noise case. Their approach to deriving a cross-coupling model from the mock/observational data results in essentially complete mitigation. Both methods show similar levels of signal loss and residual systematics (Fig. 14, bottom).
@@
Hamiltonian Monte Carlo samplers are effective at modelling and mitigating reflection and cross-coupling systematics, while also providing a statistical uncertainty estimate of the recovered foreground and EoR signals. In data containing realistic levels of noise, the systematics can be subtracted to this noise level. In essentially noise-free cases, true power spectra can be recovered following calibration, but in some cases the statistical uncertainty estimates are narrower than is likely useful. In all cases, however, sampling can provide similar performance to the current systematic mitigation techniques, while remaining resilient to signal loss, and providing statistical uncertainty measures. This can be important in both placing upper limits on, and possibly directly detecting, the 21-cm power spectrum should the noise be reduced sufficiently. Performance is very dependent on a number of factors, however, such as priors and initialisation, particularly in low-noise cases.
@@
While the signal chain systematics presented here will mostly no longer be an issue in HERA Phase II, this work was meant to demonstrate that similar systematics which lend themselves to forward modelling can likely be approached in a similar manner. In Phase II, the cables connecting the FEM to the PAM are replaced to ones with a length of 500 m, which pushes the reflections towards higher delays. The systems associated with the leaking connection point responsible for the cross-coupling are no longer in use, and so this systematic is no longer of concern. However, the introduction of Vivaldi feeds, and the lack of cages surrounding them, will likely result in over-the-air, antenna-to-antenna mutual coupling. Should this systematic be relatively time stable, and is capable of being modelled, then the techniques presented in this work may be useful for their mitigation.
@@
This work has a number of opportunities for further development. The modelling of the foregrounds and EoR contribute to the majority of the parameters in our model. Exploring alternatives to the modelling presented here could aid in reducing the impact on performance, for example only considering the broader features of these two components. For the EoR in particular, using an almost flat power spectrum in delay space could provide comparable results to the individual Fourier modes, but with more favourable sampling speed. Alternative methods of obtaining estimates for each of the components (foregrounds, EoR, and systematics) is another possible approach. While we fully intended to model each of these components, inferring them directly from mock/observational data (as is done for cross-coupling in Kern et al. (2019)) and thereafter adjusting them with a Bayesian approach could be an alternative. Good knowledge of the foreground and EoR components was also assumed beforehand, as the focus was placed on assessing the feasibility of modelling the systematics in a Bayesian framework. The effects of poor or incorrect models for the foregrounds and signal, and how this affects the mitigation of the systematics is another avenue which could be explored. Lastly, this work assumed a particular model for the sources of these systematics. It is possible that alternative models could result in an improvement in the systematics mitigation, and testing different models against one another is something Bayesian methods are ideally suited to.
@@ Acknowledgements
GM and MGS acknowledge support from the South African National Research Foundation (Grant No. 84156). This result is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 948764; P. Bull). P. Bull acknowledges support from STFC Grants ST/T000341/1 and ST/X002624/1. We acknowledge the use of the ilifu cloud computing facility – www.ilifu.ac.za, a partnership between the University of Cape Town, the University of the Western Cape, Stellenbosch University, Sol Plaatje University and the Cape Peninsula University of Technology. The ilifu facility is supported by contributions from the Inter-University Institute for Data Intensive Astronomy (IDIA – a partnership between the University of Cape Town, the University of Pretoria and the University of the Western Cape), the Computational Biology division at UCT and the Data Intensive Research Initiative of South Africa (DIRISA). We acknowledge the use of the following software: numpy (Harris et al., 2020), matplotlib (Hunter, 2007), scipy (Virtanen et al., 2020), pyuvdata (Hazelton et al., 2017), and uvtools (https://github.com/HERA-Team/uvtools).
@@
This paper makes use of simulated data from Aguirre et al. (2022). The analysis scripts used in this work are available from https://github.com/GeoffMurphy/BayesianSystematicsSampler.

### v2

@@ 6 Discussion
In this work, we present statistical modelling results when applied to the reflection and cross-coupling systematics in simulations of early HERA Phase I visibilities. These systematics serve to complicate the setting of upper limits on the EoR 21-cm signal by spreading copies of the foregrounds into regions ideally only occupied by the signal. We mimic this by adding systematics with randomised parameters to visibilities containing foregrounds and the 21-cm signal, where the initial, uncorrupted visibilities are sourced from the HERA Validation data (Aguirre et al., 2022), and the systematics are added with hera_sim. A Hamiltonian Monte Carlo sampler is used to constrain these systematics, subject to noise, and attempts are made to remove them and recover the true foreground + EoR signal. We consider both an autocorrelation case, as well as a short, 14.7 m east-west cross-correlation.
We test our model with and without an included signal model, and find that results are similar, though the signal-free model is preferred for its gains in sampling speed. At best, including a signal model provides slightly narrower uncertainty estimates in some cases, but this can lead to prohibitive runtimes, especially when noise is low or negligible.
The primary limitation of our method is the confusion between systematics, where nearly overlapping features makes it difficult for the model to accurately constrain the parameters, leading to residual systematics. This is most evident in cases where the noise is negligible, and more so for subreflection modelling in the autocorrelations, where multiple systematics peaks can occupy a small delay range. This is also a limitation of Kern et al. (2019)’s method, the currently employed calibrator in the HERA pipeline.
For autocorrelations, the twenty subreflections spread between $\sim 200-1000$ ns, and the coupling peaks at $\sim 900-1300$ ns, likely makes it difficult for the sampler to disentangle the power contributions from the multiple overlapping components, resulting in significant residual systematics in visibilities with low noise (Fig. 9). To reduce this effect would likely require very specific priors on the systematics, something which is not always feasible given slight variations in array geometry, noise, RFI, and other complications in the observations.
For cross-correlations, only ten cross-coupling spectra are modelled. Furthermore, these are high power relative to the foreground peaks, and are relatively unconfused, resulting in only slight residual systematics in low noise visibilities (Fig. 14). However, we only consider these unrealistic, essentially noise-free scenarios in order to evaluate signal loss.
With higher noise levels, the model is less confusion-limited, as a significant portion of the systematics are either equivalent in power to, or below, the noise level. This leaves many of these systematics poorly constrained, which in this case is a non-issue, as this still allows for mitigation down to this high noise level (although further averaging/integration could cause them to emerge again). For low noise, these systematics require good constraints in order to mitigate them to any significant degree. For stronger, more isolated systematics like high-amplitude cable reflections, this is relatively easy to achieve, but for the more numerous, closely-spaced systematics (e.g. subreflections and coupling), finding good constraints is more difficult.
The difficulty in modelling the systematics in the low noise scenario is further exacerbated by the reduction in computational speed as a result of this noise level, which hinders quick tests of priors and sampler hyperparameter settings, as well as the model’s ability to efficiently explore the posterior space.
@@
In autocorrelation power spectra, cable reflections, cable subreflections, and cross-coupling are of concern. These are analysed in Section 5.1. When considering a single 10.73 second time integration, we are able to constrain a number of high-amplitude reflection systematics (Fig. 7), and can mitigate down to the (relatively high) noise level. We apply this model to thirty time integrations separated by 21.46 seconds, and the recovered power spectra are incoherently averaged across time (Fig. 8). This allows for a reduction in the noise level following systematics mitigation. These reflection systematics are modelled in the autocorrelations and applied to their lower-amplitude counterparts in the corresponding cross-correlations. For the same time integration, however, reflections in cross-correlations are overwhelmed by the noise by around four orders of magnitude. As such, there is not much to be gained from calibrating out reflections in the cross-coupling in this high noise case.
@@
When the model is run on visibilities with very low noise levels, however, the systematics in the autocorrelations can be mitigated such that the recovered power spectrum is, at worst, around three orders of magnitude above the true power spectrum (Fig. 9). While this is extreme, using these results to subtract the reflections from the corresponding low noise cross-correlation results in essentially complete recovery of the true power spectrum in Fig. 10 (top). In this low noise case, there is minimal signal loss or residual systematics (Fig. 10, bottom). Furthermore, as a result of the low noise, the constraints placed on the systematics parameters become very strong, resulting in narrow statistical uncertainty estimates.
In both the high noise and low noise cases, the non-linear optimisation method of Kern et al. (2019) performs similarly to our sampler when mitigating reflections. In the former case, systematics are mitigated to the noise level, and the noise can be further reduced by incoherently time averaging. In the latter case there are some residual systematics in the autocorrelation power spectrum, but there is very good recovery of the cross-correlation with minimal signal loss. It is in the cases where noise is more realistic that the value of a Bayesian approach is more evident, as it provides some measure of statistical uncertainty in the results. For lower noise levels, these estimates become narrower than is likely useful.
@@
For a single time integration, the strongest cross-coupling peaks are about as strong as the noise, eliminating any possibility of meaningful constraints being placed the systematics. Instead, some amount of noise reduction was required. Assuming a simple time integration increase from $\Delta t=10.73\,\rm{s}$ to $1073\,\rm{s}$ reduces the noise enough to allow constraints to be placed on the highest amplitude cross-coupling peaks in Fig. 11, leading to their mitigation down to the noise level. This noise reduction is not representative of the methods used in HERA, but we take this approach to evaluate the model in the relatively high noise regime.
Once again mitigating the coupling systematics in thirty integrations separated by 21.46 seconds and incoherently averaging results in a further reduction of the noise level (Fig. 12). As with other results, our model performs comparably to the SVD method of Kern et al. (2019), within our uncertainty estimate. This required a different implementation of the sampler. All other results were found by sampling for a single visibility at a time. This result required a joint modelling approach, wherein the likelihood function had thirty terms, one for each time integration. All of these terms shared the same coupling systematics parameters as they are time-stable. By simultaneously modelling on additional data, the estimates on the parameters improved enough to allow for not only mitigation to the noise level, but further reduction of the noise after averaging, suggesting that the level of residual systematics in the recovered data was not significant.
When we evaluate a very low noise case in order to evaluate the model in an ideal scenario, i.e. where the noise power spectrum is around $10\%$ that of the EoR power spectrum, we are able to mitigate the cross-coupling spectrum down to, at worst, an order of magnitude above the true power spectrum, with there being recovery at a fair number of delays (Fig. 14, top).
@@
In this low noise modelling of the coupling systematic, Kern et al. (2019)’s SVD method performs similarly to our result. Their approach to deriving a cross-coupling model from the mock/observational data results in essentially complete mitigation. Both methods show similar levels of signal loss and residual systematics in terms of absolute dex (Fig. 14, bottom), although our model is more prone to undersubtraction, while Kern et al. (2019)’s to oversubtraction.
@@
Hamiltonian Monte Carlo samplers are effective at modelling and mitigating reflection and cross-coupling systematics, while also providing a statistical uncertainty estimate of the recovered foreground and EoR signals. By modelling corrupted visibilities containing realistic levels of noise, the systematics can be subtracted to this noise level, and incoherently averaging the recovered spectra can further supress this noise. In essentially noise-free cases, true cross-correlation power spectra can be recovered with minimal signal loss and residual systematics, but the statistical uncertainty estimates are often narrower than is likely useful. In most cases, however, sampling can provide similar performance to the current systematic mitigation techniques, while remaining resilient to signal loss and providing statistical uncertainty measures. This can be important in both placing upper limits on, and possibly directly detecting, the early-Universe 21-cm power spectrum should the noise be reduced sufficiently. Performance is very dependent on a number of factors, however, such as priors and initialisation, particularly in low noise cases.
For HERA, the signal-chain systematics subtraction takes place after the majority of the analyses, namely redundant calibration, absolute calibration, RFI flagging, in-painting, etc. Only coherent time-averaging takes place after systematics removal, and thereafter power spectra are formed (Aguirre et al., 2022). As such, implementing a method similar to what is presented here would be fairly straightforward. Following systematics removal, instead of a single visibility for each time and each antenna, there would be however many sample visibilities. One option is to pass all of these visibilities through the pipeline and thousands of power spectra could be formed for each time and each antenna, from which uncertainties could be estimated. A less computationally expensive alternative is to form the uncertainties on the recovered visibilities, as was done in this work, and propagate those through the pipeline. For example, one could replace the observed visibilities with the upper and lower bounds of the percentile regions, effectively only doubling the amount of data, rather than increasing it by $\mathcal{O}(10^{5})$ or more.
@@
While the signal chain systematics presented here will mostly no longer be an issue in HERA Phase II, this work was meant to demonstrate that similar systematics which lend themselves to forward modelling can likely be mitigated in a similar manner. In Phase II, the cables connecting the FEM to the PAM are replaced by ones with a length of 500 m, which pushes the reflections towards higher delays. The systems associated with the leaking connection point responsible for the cross-coupling are no longer in use, and so this systematic is no longer of concern. However, the introduction of Vivaldi feeds, and the lack of cages surrounding them, will likely result in over-the-air, antenna-to-antenna mutual coupling. Should this systematic be relatively time stable, and be capable of forward modelling, then the techniques presented in this work may be useful for their mitigation.
This work has a number of opportunities for further development. For the realistic noise cases, reflection modelling in the autocorrelations takes approximately 60 minutes for each individual time snapshot, given the large number of parameters and the use of the ADVI initialiser. This time includes model initialisation for two chains, a total of 8,000 tuning steps, and 10,000 sampling steps. Coupling modelling in a single cross-correlation, and for a single integration, is only of the order 10 minutes due to the reduced number of parameters and the use of a less computationally expensive initialiser. However, multiple integrations are included in the likelihood function for better results. Jointly modelling thirty integrations with two chains for a total of 10,000 tuning steps and 10,000 sampling steps takes approximately 3.5 hours.
For visibilities with negligible noise, run times extend into multiple hours for an individual time snapshot with similar settings. While levels of noise this low are not currently obtainable, for future data this would be an unreasonably high runtime. This is mainly due to the assumed noise level itself, where less noise typically results in slower sampling.
@@
While we fully intended to model each of these components, inferring them directly from mock/observational data, as is done for cross-coupling in Kern et al. (2019), and thereafter adjusting them with a Bayesian approach could be an alternative. Good knowledge of the foreground components was also assumed beforehand, as the focus was placed on assessing the feasibility of modelling the systematics in a Bayesian framework. The effects of poor or incorrect models for the foregrounds and signal, and how this affects the mitigation of the systematics is another avenue which could be explored. Lastly, this work assumed a particular model for the sources of these systematics. It is possible that alternative models could result in an improvement in the systematics mitigation, and testing different models against one another is something Bayesian methods are ideally suited to.
@@ Acknowledgements
Analysis, writing, and editing was carried out by GGM, PB, and MGS. Further analysis of results was carried out by PK and JSD. The remaining authors, listed in alphabetical order, are members of the HERA Builder’s List, which recognises their direct contributions to the construction, commissioning, and operation of the telescope and analysis pipeline development that are a necessary basis of the HERA-specific data and specifications used within this paper.
We thank the anonymous reviewer for their feedback, which has resulted in significant improvement to the paper.
GGM and MGS acknowledge support from the South African National Research Foundation (Grant No. 84156). This result is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 948764; P. Bull). P. Bull acknowledges support from STFC Grants ST/T000341/1 and ST/X002624/1. We acknowledge the use of the ilifu cloud computing facility – www.ilifu.ac.za, a partnership between the University of Cape Town, the University of the Western Cape, Stellenbosch University, Sol Plaatje University and the Cape Peninsula University of Technology. The ilifu facility is supported by contributions from the Inter-University Institute for Data Intensive Astronomy (IDIA – a partnership between the University of Cape Town, the University of Pretoria and the University of the Western Cape), the Computational Biology division at UCT and the Data Intensive Research Initiative of South Africa (DIRISA). We acknowledge the use of the following software: numpy (Harris et al., 2020), matplotlib (Hunter, 2007), scipy (Virtanen et al., 2020), pyuvdata (Hazelton et al., 2017), and uvtools (https://github.com/HERA-Team/uvtools).
@@
This paper makes use of simulated data from Aguirre et al. (2022). The analysis scripts used in this work are available from https://github.com/GeoffMurphy/HMCSystematicsSampler.